### v1

Abstract
One of today’s most significant and transformative technologies is the rapidly developing field of artificial intelligence (AI). Defined as a computer system that simulates human cognitive processes, AI is present in many aspects of our daily lives, from the self-driving cars on the road to the virtual assistants in our smartphones. The term ”black box” is often used to describe artificial intelligence (AI) because some AI systems are so complex and opaque. With millions of parameters and layers, these systems—deep neural networks in particular—make it difficult for humans to comprehend how they make judgments. Even if AI is capable of producing correct findings, questions of accountability, prejudice, and justice are raised by the opaqueness of its decision-making process. AI has a lot of potential, but it also comes with a lot of difficulties and moral dilemmas. In the context of explainable artificial intelligence (XAI), trust is crucial as it ensures that AI systems behave consistently, fairly, and ethically. In the present article, we explore XAI, reliable XAI, and several practical uses for reliable XAI. Once more, we go over the three main components—transparency, explainability, and trustworthiness of XAI—that we determined are pertinent in this situation. We present an overview of recent scientific studies that employ trustworthy XAI in relation to these fundamental components, as well as an analysis of how trustworthy XAI is applied in various application fields. In the end, trustworthiness is crucial for establishing and maintaining trust between humans and AI systems, facilitating the integration of AI systems into various applications and domains for the benefit of society.
keywords:
Artificial Intelligence(AI), Trustworthy XAI, Explainable Artificial Intelligence (XAI), Healthcare, Autonomous Vehicles
1 Introduction
Philosophers who sought to characterize human thought as the mechanical manipulation of symbols laid the foundations of modern-day artificial intelligence. These efforts resulted in the invention of the programmable digital computer [1] in the 1940s. Though it is now lost, Alan Turing may have published the first article on the subject of artificial intelligence in 1941, indicating that he was at least thinking about the concept at that time. The public was first introduced to Turing’s notion of the Turing test in his seminal work ”Computing Machinery and Intelligence” from 1950 [2]. In it, Turing cast doubt on the viability of building thinking machines. The term artificial intelligence (AI) was first used in 1950 [3] but its general adoption and use in healthcare has been hindered by many shortcomings of the original models. The introduction of deep learning in the early 2000s eliminated many of these limitations. We are entering a new era of technology where AI can be used in clinical practice through risk assessment models that improve workflow efficiency and diagnostic accuracy. AI systems can now analyze complex algorithms and learn on their own. The performance of AI systems has improved significantly in recent years. These new models expand on their capabilities to include text-image synthesis based on nearly any prompt, whereas previous systems focused primarily on generating facial images.
Figure 1: Domains throughout which artificial intelligence finds applications. [4]
The applicability and potential of artificial intelligence (AI) to transform business is already evident in the wide range of areas in which it is applied: In the field of natural language processing (NLP) [5], artificial intelligence (AI) makes it easier for computers to understand and generate human language, enabling tasks such as sentiment analysis, machine translation, and spam filtering. Furthermore, computer vision [6] enables computers to comprehend visual information, promoting developments in fields like object identification, facial recognition, and self-driving automobiles. Computers can now learn from data thanks to machine learning (ML), which has applications in fraud detection, recommendation systems, predictive analytics, and other fields. Robotics [7] is a branch of artificial intelligence that deals with the design, development, and use of machines. These machines are used in a variety of industries, including space exploration, manufacturing, healthcare [8], [9], and many more. Additionally, the incorporation of AI into business intelligence (BI) [10] signals better data gathering, processing, and display, which promotes data-driven decision-making and increases efficiency. AI offers advances in patient outcomes and medical developments in the healthcare industry [11],[12], [13] by assisting in illness diagnosis, therapy development, and tailored care. AI’s promise in education is evident in its ability to personalize instruction, engage students, and automate administrative tasks—all of which open the door to more customized learning opportunities. Through data-driven tactics, artificial intelligence (AI) in agriculture maximizes crop productivity, lowers expenses, and assures environmental sustainability. Similar to this, AI in manufacturing uses task automation and process optimization to increase productivity, efficiency, and quality. AI’s influence is felt not just in these domains but also in banking, retail, energy, transportation [14], handwriting detection [15], and government, where it is transforming operations, improving services, and reshaping global industry landscapes. Figure 1 illustrates the wide range of industries in which artificial intelligence finds widespread use. These industries include retail, security, healthcare, e-commerce, manufacturing, finance, transportation and logistics, and home furnishings. These apps depend on reasonably developed AI technologies including ML, NLP, and computer vision.
1.1 Third Wave of Artificial Intelligence (3AI)
The majority of commercial AI technology today is classified as ”narrow AI,” meaning that it consists of extremely specialized systems that excel at a limited number of clearly defined jobs and nothing else. Even the most amazing autonomous cars need a combination of limited artificial intelligence algorithms. The reliance of modern AI on enormous training data sets is another drawback. For example, a three-year-old child can recognize cats from only a few instances, while a standard machine learning algorithm would need to be given tens of thousands of cat photographs before it can recognize them with any degree of precision. The idea of ”Third Wave AI” emerges because artificial intelligence has to become more humanlike in a number of ways in order to overcome these constraints and realize its full potential.
Figure 2: Past, Present, and Future of AI waves. [16]
According to the Defense Advanced Research Projects Agency (DARPA) [17], third-wave AI systems will be able to understand context, apply that contextual awareness with common sense, and adapt to changing conditions. This will make it possible for AI systems and human users to connect in a more organic and intuitive way [17]. One of the few DARPA projects underway, XAI is anticipated to pave the way for ”third-wave AI systems,” in which computers comprehend the context and surroundings in which they function and gradually develop the underlying explanatory models necessary to describe occurrences in the actual world.
  • First Wave AI was centered on rules, logic, and constructed knowledge.
  • Big data, statistical learning, and probabilistic techniques were introduced by Second Wave AI.
  • The goal of third-wave AI is to develop common sense and contextual adaption skills.
According to Tractica [16], the worldwide market for artificial intelligence software is predicted to generate revenues of 118.6 billion by 2025, up from around 9.5 billion US dollars in 2018. The focus of this data is on developing AI systems that not only perform tasks accurately but also provide explanations of their decisions and actions in ways that humans can understand.
The third wave of AI refers to the evolution of AI technologies beyond traditional machine learning methods towards more advanced approaches that integrate reasoning, context awareness, and human-like understanding. This wave aims to develop AI systems capable of understanding and interacting with the world in more nuanced and sophisticated ways, often drawing inspiration from cognitive science and neuroscience.
Explainable AI (XAI), on the other hand, focuses on improving the transparency and interpretability of AI systems, especially machine learning models, so that humans can understand and trust the decisions made by these systems. XAI techniques aim to provide explanations of AI predictions and actions, enabling users, including developers, regulators, and end users, to understand the underlying considerations and factors that influence AI outcomes.
The relationship between the third wave of AI and XAI lies in their shared goal of creating AI systems that are not only capable of making accurate predictions or decisions but also capable of providing meaningful explanations for those predictions or decisions. By incorporating XAI techniques into the development of third-wave AI systems, developers can ensure that these systems are not only advanced in their capabilities but also transparent and interpretable, fostering trust and acceptance among users.
1.2 Concept of Explainable AI
Artificial Intelligence is often criticized for being hard to explain. Many opponents argue that it is hard to trust the results of an AI if one does not know how it can arrive at a certain result or conclusion. This issue becomes particularly problematic when AI-based programs and systems are unable to accomplish their goals. Developing explainability is necessary to boost public confidence in the computational execution. Figure 3 shows the AI in use today and the AI expected to be used tomorrow. If we are to hold the system accountable, we must take steps to address and minimize input inefficiencies [18].
Artificial intelligence is known to suffer from the ”black box” syndrome due to a lack of understanding of how the system works. This has ramifications for human confidentiality, arbitrary discrimination, obfuscation, and legitimacy. There are often underlying prejudices and tendencies present along with this lack of openness. By enhancing users’ comprehension of how AI-powered systems reason, XAI seeks to improve their performance. The goal of Transparent AI is to make artificial intelligence (AI) safer and more accessible than ever before [18]. Therefore, each of the TAI’s diverse features has to be considered separately, and their many facets need to be talked about.
AI’s use of machine learning (ML) techniques falls into two categories: white box and black box [19]. The results of a white box model are understandable to a subject matter expert. In contrast, a black box model is very difficult to explain and can be confusing even to domain experts [20]. The three criteria of interpretability, explainability, and transparency are adhered to by the XAI algorithm [21]. A model is transparent when the underlying principles of the machine learning model and decision-making can be understood ”when the process of extracting model parameters from training data and generating labels from test data is described and justified by the designers of the approach. ” Because communicating in a way that people can understand is called interpretability [22]. Currently, there is no universally accepted definition of the concept of explainability, but its importance is recognized. An alternative is ”a collection of interpretable features of a domain that can help generate a decision (such as classification or regression) about a given example.” If an algorithm follows these guidelines, it provides a basis for recording and validating the decision, improving the algorithm, and discovering new information.
Figure 3: Explainable Artificial Intelligence (XAI): A look at AI now and tomorrow. [18]
From an explainability perspective, academics are actively studying intelligent systems, which is a key topic. Formal compliance may occasionally depend on the system’s understanding. Many black box algorithms, as shown in Figure 4, present a trade-off between incredible learning performance (accuracy) and provable AI.
Figure 4: Performance of artificial intelligence vs explainability. [23]
1.3 Classification Tree of XAI
XAI techniques are divided into two categories: transparent and post-hoc methods. A transparent approach is one that represents the model’s capabilities and decision-making process in an easy-to-understand way [24]. Transparent models include Bayesian approaches, decision trees, linear regression, and fuzzy inference systems. Transparent approaches can be useful when the internal feature correlations are highly complex or linear. A comprehensive classification of different XAI methods and approaches related to different types of data is shown in Figure LABEL:Classification [lertvittaakumjorn2019human].
Figure 5: XAI categorization according on data type [24]
Posterior approaches are useful for interpreting the complexity of a model, especially when there are nonlinear relationships or high data complexity. When a model does not follow a direct relationship between data and features, posterior techniques can be an effective tool to explain what the model has learned [24]. Inference using local feature weights is provided by transparent methods such as Bayesian classifiers, support vector machines, logistic regression, and K-nearest neighbors. This model category meets three properties: simulability, decomposability, and algorithmic transparency [24].
1.4 Definition of Transparency in Artificial Intelligence
In explainable artificial intelligence (XAI), transparency refers to the ability of an AI system to provide understandable justification for its decisions and actions [25]. In fact, transparency is one of the key components of explainable artificial intelligence (XAI). In many real-world applications, particularly those with large social implications, deciphering the reasoning behind an AI system’s choice is just as crucial as the decision itself. Assuring transparency in XAI helps prevent AI systems from being viewed as ”black boxes” and instead as instruments that enable users to get insightful and informative responses [26].
The general transparency of AI systems is further enhanced by the insights offered by XAI approaches. Users have the capacity to scrutinize the decision-making procedures, detect any partialities, and evaluate the dependability and equity of the model’s results [27]. Transparent systems are essential for guaranteeing accountability and ethical concerns in fields where the implications of AI choices might have large repercussions, such as healthcare, finance, and autonomous cars. XAI approaches enable users to find patterns, comprehend relationships, and discover any flaws or biases by providing insightful information about the inner workings of AI models [27]. Stakeholders are better equipped to make judgments, confirm that the model’s predictions are accurate, and take necessary action as a result of the improved openness.
The examination of the ethical criteria revealed a correlation between explainability and transparency and a number of other quality needs. The nine quality standards pertaining to explainability and openness are shown in Figure 6.
Figure 6: Qualitative standards of explainability and openness + Assistance; – Disagreements. [28]
The growth of AI systems’ explainability and transparency is facilitated by their understandability. When discussing the significance of understandability, the transparency guidelines addressed three points: 1) ensuring that people comprehend the AI system’s behavior and the methods for using it (O5, O12); 2) communicating in an intelligible manner the locations, purposes, and methods of AI use (O15); and 3) making sure people comprehend the distinction between real AI decisions and those that AI merely assists in making (O2) [28]. Thus, by guaranteeing that people are informed about the use of AI in a straightforward and comprehensive manner, understandability promotes explainability and transparency. The necessity of tracking the decisions made by AI systems is highlighted by traceability in transparency requirements (O2, O12) [28]. In order to ensure openness, Organization O12 also noted how crucial it is to track the data utilized in AI decision-making.
1.5 Transparency Vs Explainability in AI
Transparency and explainability can be compared [29]. When an AI is transparent, the ”basic elements of data and decisions must be available for inspection during and after AI use,” according to McLarney et al. [30]. Transparency exists when a user can observe how decisions are made or have access to their data. Explainability, on the other hand, is about understanding why AI succeeds or fails and revealing how it draws on the knowledge and decision-making processes of the people it will affect. It provides a rational explanation for the AI’s actions. Users need to be able to understand what data is being collected, how the AI program processes it, and how it produces trustworthy outcomes for each individual affected. This straightforward explanation ignores the challenges we confront in simplifying ”black box” algorithms, the context that is lost, and the accuracy needed when giving consumers clear explanations. The question thus becomes, is minimal explainability preferable to nothing? [30]. Other important factors to consider include the belief that explanations can adequately account for the dynamic nature of the rich information ecosystem and the appropriateness of dealing with anomalies.
Figure 7: Output from the Bing search engine’s conversation feature explaining a failure. a partial screenshot taken using an Android smartphone on March 2, 2023. [17]
Interestingly, while certain AI algorithms analyze data automatically, an increasing number of AI systems are designed to explain how their algorithms work and the reasoning behind certain decisions [17]. For example, the Bing search engine’s conversation mode offers brief explanations of how it works (Fig. 7). End users may occasionally find these explanations enough, but occasionally they may get confused about how an AI arrived at a specific conclusion or behaved in a specific way. It is impractical to expect people to become more computer literate when they are more perplexed by the explanation that is provided [17]. Rather, we need to either enhance the AI system or the justification.
1.6 Definition of Trustworthiness in Artificial Intelligence
Artificial intelligence (AI) systems that incorporate trustworthiness must take a multifaceted approach that takes organizational, ethical, and technical factors into account. Establishing criteria for assessing trustworthiness is the first step in this process. These measures should include accountability, security, privacy, openness, fairness, and ethical compliance. Foundational elements include transparent algorithms that provide intelligible justifications for AI-driven judgments and high-quality, impartial data. Strong security measures and privacy-preserving strategies protect sensitive data and fight off online attacks. Responsible AI usage is encouraged through the establishment of accountability systems and adherence to moral principles and governance structures. User-centric design, ongoing observation, and training guarantee that AI systems satisfy users’ demands while developing over time to retain their credibility. Organizations may create trustworthy, transparent, equitable, and ethical systems that inspire confidence in both users and stakeholders by incorporating these principles into all phases of the AI lifecycle.
Figure 8: The three primary elements of a reliable AI. [31]
The three elements depicted in Figure 8 - algorithmic ethics, data ethics, and practice ethics - come together to form trustworthy AI. These elements offer an abstraction level for ethical issues that is data-centric [31]. Numerous unresolved problems arise when trying to address ethical concerns with AI systems. In the research paper [31], authors describe the requirements for Trustworthy AI where the mentioned:
  1. 1)
    Human agency and oversight: AI systems must uphold the values of human freedom. AI systems must promote user agency, uphold fundamental rights, and enable human oversight to realize a democratic and equal society.
  2. 2)
    Security and technical robustness: Security and technical robustness lead to the prevention of harm. For an AI system to function reliably while minimizing harm, risks must be taken into account during development. This must cover any changes in the working environment as well as possible attacks on the system by adversaries.
  3. 3)
    Data protection and data governance: As a fundamental right that is particularly affected by the extensive data collection required by AI systems, data protection is also closely related to preventing harm. Preventing privacy harm also requires data governance that addresses the quality and integrity of the data used, its relevance, access protocols, and the ability to process it in a privacy-preserving way.
  4. 4)
    Transparency: Explainability and transparency are closely related requirements. The aim is to make all relevant aspects of an AI system transparent, including data, technology, and business models. In the era of ubiquitous computing, transparency is essential to support large-scale data collection and its benefits to consumers.
  5. 5)
    Diversity, non-discrimination, and fairness: Achieving trustworthy AI requires enabling inclusivity and diversity throughout AI systems. This is important not only to consider and involve all affected parties but also to ensure fair access and treatment. Fairness and this need go hand in hand.
  6. 6)
    Social and environmental welfare: In the spirit of justice and harm prevention, the environment and the broader community should be considered as stakeholders. Research into AI solutions to address global challenges should be promoted, making AI systems more environmentally friendly and sustainable. AI systems need to benefit everyone, including those who will come after us.
  7. 7)
    Accountability: The idea of fairness and the need for accountability go hand in hand: we need mechanisms to ensure accountability and accountability for AI systems and the results they produce, both during and after their development, application, and deployment.
1.7 An Overview of Necessities for Reliable AI
The conditions for reliable AI are still unclear and are addressed incongruously by many institutions and groups, despite contentious social debates over it. The principles of Fairness, Responsibility, Accuracy, Verifiability, and Accountability in Machine Learning (FAT-ML) include accountability, explainability, verifiability, and fairness at a global level [32]. Among the numerous needs under review, explainability, fairness, privacy, and robustness will all be covered in this study (Table 1).
Table 1: Conditions necessary for trustworthy artificial intelligence (AI)
Concept Description
Explainability To help consumers comprehend, the method by which the AI model generates its output might be demonstrated.
Fairness Regardless of certain protected variables, the AI model’s output can be shown.
Privacy It is feasible to prevent issues with personal data that might arise while the AI is being developed.
Robustness The AI model can fend against outside threats while continuing to operate correctly.
2 Trustworthy XAI Vs AI
A paradigm change in the field of artificial intelligence (AI) has been brought about with the introduction of Trustworthy Explainable AI (XAI). Oftentimes, conventional AI systems operate as opaque black boxes, making it challenging for users to comprehend the decision-making process. On the other hand, Trustworthy XAI tackles significant concerns surrounding the adoption of AI by focusing on accountability, interpretability, and transparency. Reputable XAI seeks to build consumers’ confidence by providing clear justifications for its choices. Users are able to evaluate the fairness and dependability of AI-driven results because to this transparency. While typical AI systems are capable of producing precise forecasts or suggestions, they do not have the openness required to establish credibility.
The way that reliable XAI and traditional AI make decisions is what sets them apart from one another. Although AI systems have the potential to produce precise forecasts or suggestions, Reputable XAI highlights the necessity of providing justifications that clarify the processes that lead to these results. Users may evaluate the fairness and dependability of AI-driven outcomes thanks to the clear justifications that trustworthy XAI systems give for their decisions.
Furthermore, trustworthy XAI takes into account more ethical factors than just explanations. Incorporating the Fairness, Accountability, and Transparency (FAT) principles into AI development procedures, guarantees that AI systems abide by moral and legal requirements. Trustworthy XAI seeks to reduce biases, discrimination, and other possible negative effects of AI technology by placing a high priority on ethical norms. An approach to AI development known as trustworthy AI places a high value on user safety and openness. Since no model is flawless, trustworthy AI developers take care to explain to clients and the wider public how the technology was developed, its intended applications, and its limitations.
Table 2: Seven Requirements to Meet in Order to Develop Reliable AI
Principles Explanation Rights GDPR Ref
Human Authority and Supervision Artificial intelligence technology ought to uphold human agency and basic rights, instead of limiting or impeding human autonomy. The right to get human assistance Recital 71, Art 22
Robustness and Safety Systems must be dependable, safe, robust enough to tolerate mistakes or inconsistencies, and capable of deviating from a totally automated decision Art 22
Data Governance and Privacy Individuals should be in total control of the information that is about them, and information about them should not be used against them Notification and information access rights regarding the logic used in automated processes Art 13, 14, and 15
Transparency Systems using artificial intelligence ought to be transparent and traceable The right to get clarification Recital 71
Diversity and Fairness AI systems have to provide accessibility and take into account the whole spectrum of human capacities, requirements, and standards Right to not have decisions made only by machines Art 22
Environmental and Social Well-Being AI should be utilized to promote social change, accountability, and environmental sustainability Accurate knowledge regarding the importance and possible consequences of making decisions exclusively through automation Art 13, 14, and 15
Accountability Establishing procedures to guarantee that AI systems and their outcomes are held accountable is essential Right to be informed when decisions are made only by machines Art 13, 14
3 Applications of Trustworthy XAI
Authentic Explainable Artificial Intelligence (XAI) has numerous uses in sectors where accountability, interpretability, and transparency are essential. XAI can provide an explanation for a diagnosis or therapy recommendation in medical diagnosis and recommendation systems. Financial institutions can employ XAI for risk assessment, fraud detection, and credit scoring. XAI can help attorneys with contract analysis, lawsuit prediction, and legal research. In autonomous vehicles, XAI plays a significant role in providing context for the decisions made by the AI systems, particularly in high-stakes scenarios such as accidents or unanticipated roadside incidents. XAI can be applied to process optimization, predictive maintenance, and quality control in manufacturing settings. By offering justifications for automated responses or suggestions in chatbots and virtual assistants, XAI can improve customer service. By providing an explanation for the recommendations and assessments made by adaptive learning systems, XAI can help with individualized learning. By providing an explanation for the recommendations and assessments made by adaptive learning systems, XAI can help with individualized learning. We shall concentrate on a few particular applications in this section and go into detail about them.
3.1 Application of Trustworthy XAI in Medical Science
The field of artificial intelligence (AI) is rapidly growing on a global scale. The potential uses of artificial intelligence in healthcare are a hot topic for research [33]. There are many opportunities to use AI technology in the healthcare sector, where people’s lives and well-being are in danger, because of its essential relevance and the enormous quantity of digital medical data that have been gathered [34]. Artificial intelligence (AI) has made it possible to accomplish tasks quickly that were before unfeasible for traditional technologies. Trustworthy AI is a huge concern these days. Since incidents involving AI-powered chatbots such as Tay and Iruda [35], there has been an increase in interest in the topic of whether an AI’s judgment and decision-making system is reliable. The credibility of AI in the medical and healthcare areas requires more investigation. Clinical decision support systems (CDSS) in the medical field use AI technology to aid with important medical duties such as diagnosis and therapy planning [36]. Misuse can have serious consequences in areas where lives are at stake, even if the scope of use is limited to assisting healthcare practitioners. False alarms, for example, which occur often in scenarios involving urgent patients, may exhaust medical personnel.
The study [37] adds significantly to the field of medical skin lesion diagnostics in a number of ways. Before anything else, it adapts an existing eXplainable Artificial Intelligence (XAI) technique to increase user confidence and trust in AI decision-making systems. This modification entails describing an AI model that is skilled at differentiating between different kinds of skin lesions. Synthetic exemplar and counter-exemplar images are used to create explanations that illustrate the important characteristics that influence classification choices. This research [37] is based on training a deep learning classifier with the ISIC 2019 dataset using the ResNet architecture. This enables practitioners to use the explanations offered to reason effectively. All things considered, the study’s original contributions are found in the way it refined and assessed the XAI technique in an actual medical setting, examined the latent space and carried out an extensive user study to gauge the efficacy of the explanations, especially among subject matter experts.
This research paper [38] addresses the challenge of recognizing brain tumors in MRI images by merging two powerful algorithms: fuzzy C-means (the FCM method) and Artificial Neural Network (ANN). The authors want to increase the segmentation process’s interpretability as well as the accuracy of tumor identification by merging these techniques. Their main objective is to improve medical decision-support system tools so that physicians can diagnose brain tumors more accurately. This discovery has two main benefits: first, it improves the ability to identify brain cancers in medical imaging with more accuracy, which is important for early diagnosis and treatment. Second, the researchers make the decisions made by their models more transparent and intelligible to patients and medical experts by integrating explainable AI principles into the segmentation process. In the end, this improved interpretability raises the general level of reliability and acceptance of AI-driven systems for medical picture segmentation in clinical settings.
In another research [39], the field of computational pathology—which uses machine learning and artificial intelligence (AI) to diagnose whole slide images (WSIs)—is discussed. Because artificial intelligence is opaque, there are doubts regarding its reliability despite the technology’s potential to improve efficiency and accuracy. The paper suggests employing explainable AI (xAI) techniques, which can shed light on the choices made by AI algorithms, to allay these worries. Computational pathology systems become more clear and dependable with the addition of xAI, especially when it comes to crucial activities like pathology diagnosis. Additionally, it presents HistoMapr-Breast, a software program with xAI capabilities intended for breast core biopsies.
In a recent research [40], it is discussed how crucial it is to guarantee the correctness and resilience of AI-based systems in the healthcare industry, especially with regard to their interpretability and defense against hostile attacks. As AI systems are used in medical contexts more frequently, it is important to confirm that the predictions they are generating are based on accurate features. Numerous model interpretability and explainability techniques have been put out in an effort to overcome this. This work shows that even with strong training, adversarial attacks can affect the explainability of a model. The authors also introduce two attack classifiers: one to determine the type of attack and another to differentiate between benign and opposed inputs.
This research paper [41] studies the rising area of explainable machine learning in cardiology. It addresses challenges with the interpretability of complicated prediction models and their ramifications for key healthcare choices. This study delves into the underlying principles and tactics of explainable machine learning, providing cardiologists with a greater grasp of the approach’s strengths and limitations. The research intends to help decision-making processes in clinical settings by providing a rule of thumb for deciding between interpretable and black-box models. This will ultimately improve patient outcomes while maintaining accountability and transparency in model predictions. Figure 9 was created by training a single new decision tree based on the random forest model predictions. The global tree diagram depicts the overall operation of the random forest. Individual predictions can then be investigated by tracking the patient’s progress through the global tree. This type of explanation has the advantage of making it easy to understand both the overall operation of the model and the reasoning behind each particular forecast. Decision trees are ideal for disciplines such as cardiology because they provide rule-based reasoning similar to clinical decision criteria.
Figure 9: A random forest-based model for heart disease prediction is described, using both local and global decision trees. Because decision trees are evaluated from top to bottom, the global diagram shows that the model begins by deciding if patients’ thallium stress test results are normal. The model investigates the patient’s ST depression if the thallium stress test discloses an issue, and so on. The local graphic depicts the path a patient took to fall from the tree, showing the reasoning for each individual prognosis. The patient was less than 54.5 years old, had a maximum heart rate of higher than, and had a normal thallium stress test result, indicating a very low chance of cardiac disease [41].
Figure 10 depicts the LIME explanations for our heart failure model’s two local predictions. The authors demonstrate how these predictions are integrated as a clinical decision support tool in Epic, an electronic health record created with doctors in mind (Epic Systems Corporation, Verona, Wisconsin, USA).37 This type of explainability approach has the advantage of providing detailed explanations for the clinical factors that impact each prediction. It is noteworthy that this kind of explanation can be incorporated into an EHR, as this could enhance a black-box model’s actionability by allowing forecasts and clear explanations to be effortlessly integrated into clinical workflow.
Figure 10: Heart disease prediction explanation produced using Local Interpretable Model-Agnostic Explanations (LIME). This illustration shows how clinical decision assistance can be integrated into an Epic electronic health record by means of a local explanation utilizing the LIME algorithm. To help clinicians identify patients who are likely to be at a high risk of heart disease, probabilities are color-coded. To improve the predictability and actionability of the results for doctors, the clinical factors that are most significant to the prediction are shown on the right [41].
Reliable medical AI implementation and widespread usage need substantial research and social consensus on prerequisites such as explainability, fairness, privacy protection, and robustness [32]. Optimized requirements and standards must be satisfied in any therapeutic environment where AI is used, and these requirements and standards need to be updated often. Furthermore, laws must be put in place specifying who is in charge in the event of a medical AI-related mishap or accident—designers, researchers, medical personnel, or patients [42].
3.2 Explainability and Interpretability of Autonomous Systems
Explainability and interpretability in the context of autonomous systems refer to the ability to understand and make sense of the systems’ decisions and behaviors. Explainability refers to an autonomous system’s ability to provide clear arguments for its decisions and behaviors [43]. It is critical for increasing acceptance and confidence in AI systems, particularly in areas such as banking, healthcare, and autonomous autos. While explainability and interpretability are closely connected, interpretability places more emphasis on the capacity to comprehend the internal workings and procedures of the autonomous system [44]. An interpretable system provides users with insight into the elements and criteria considered while making decisions, allowing them to comprehend how the system came to its findings.
The research paper The research article [18] focuses on trust and dependability in autonomous systems. Autonomous systems have the potential for system operation, rapid information dissemination, massive data processing, working in hazardous environments, operating with greater resilience and tenacity than humans, and even astronomical examination [45], [46]. Following years of research and development, today’s automated technologies represent the peak of progress in computer recognition, responsive systems, user-friendly interface design, and sensing automation. According to [43], the global market for automotive intelligent hardware, operations, and innovation is expected to rise from \$1.25 billion in 2017 to \$28.5 billion by 2025. According to Intel’s research of the predicted advantages of autonomous cars, employing these advances on public roads would decrease commute time by 250 million hours annually and save over 500,000 lives in the United States alone between 2035 and 2045 [43]. Modern automobiles employ artificial intelligence (AI) for several tasks, including intelligent cruise control, automatic driving and parking, and blind-spot identification (Figure 111).
Authors [18] describe the challenges of autonomous systems, like, people sometimes tend to be overly excited about the potential of new ideas and ignore, or at least appear to be unaware of, the potential drawbacks of cutting-edge developments. Even in the early stages of robotics and autonomous system implementation, humanity preferred to put up with faulty goods and services, but they have gradually come to understand the importance of trustworthy and dependable autonomous systems [zablocki2022explainability]. Numerous examples have demonstrated how operators’ use of automation is greatly impacted by trustworthiness.
Figure 11: An automated vehicle that provides a valid and understandable rationale for the decision it made at that particular instant, acting as the archetypal example of XAI in automated driving [18].
When artificial intelligence (AI) has become prevalent in autonomous vehicle (AV) operations, user trust has been identified as a major issue that is essential to the success of these operations. Explainable artificial intelligence (XAI), which calls for the AI system to give the user explanations for every decision it makes, is a viable approach to fostering user trust for such integrated AI-based driving systems [47]. This work develops explainable Deep Learning (DL) models to improve trustworthiness in autonomous driving systems, driven by the need to improve user trust and the potential of innovative XAI technology in addressing such requirements. The main concept of this [47] research is to frame the decision-making process of autonomous vehicles (AVs) as an image-captioning task, generating textual descriptions of driving scenarios to serve as understandable explanations for humans. The proposed multi-modal deep learning architecture (Shown in Figure 12), based on Transformers, effectively models the correlation between images and language, generating meaningful descriptions and driving actions. Its contributions lie in formulating the traditional AV decision process for explainability, developing a fully Transformer-based model for generating descriptions and actions, and demonstrating superior performance over baseline models. The outcome is a model that enhances user trust, provides insights for AV developers, and offers superior interpretability through its attention mechanisms and end-to-end goal induction.
Figure 12: The Transformers-based multi-modal deep learning architecture that is being suggested [47]
This research [48] aims to investigate the integration of explainable artificial intelligence (XAI) into autonomous vehicular systems to improve transparency and human trust. It delves into the functioning of multiple inner vehicle modules, emphasizing the importance of understanding the vehicle’s decision-making processes for user credibility and reliability. The main contribution lies in introducing XAI to the domain of autonomous vehicles, showcasing its role in fostering trust, and highlighting advancements through comparative analysis. The output comprises the creation of visual explanatory techniques and an intrusion detection classifier, which show considerable advances over previous work in terms of transparency and safety in autonomous transportation systems.
3.3 Applications of XAI for Operations in the Industry
The process industry is a subset of businesses that manufacture items from raw materials (not components) using formulae or recipes. Given the magnitude and dynamic nature of operations in the process sector, it becomes evident that the next great step ahead will be the capacity for people and AI systems to collaborate to ensure production stability and dependability [49]. AI systems must successfully inform the individuals who share the ecosystem about their objectives, intentions, and findings as the first step toward collaboration. In the future, people will work ”with” automation rather than ”around” it, thanks in part to the systematic approach to XAI.
This research [50] focuses on Explainable Artificial Intelligence (XAI) applications in the process industry. The research argues that current AI models are not transparent enough for process industry applications, and highlights the need for XAI models that can be understood by human experts. The main contribution is outlining the challenges and research needs for XAI in the process industry. The outcome is to develop XAI models that are safe, reliable, and meet the needs of human users in the process industry.
Table 3: Examples of AI applications in process industry operations, including pertinent data, users, and procedures. (RNN = Recurrent Neural Network; KNN = K-Nearest Neighbor; ANN = Artificial Neural Network; SVM = Support Vector Machine; SVR = Support Vector Regression; RF = Random Forest; IF = Isolation Forest) [50]
Reference Relevant Data End Users Application AI Methods
[51], [52], [53] Process signals Operator, Process Engineer, Automation engineer Process monitoring RNN, KNN
[54], [55], [56] Process signals, Alarms, Vibration Process engineer, Automation engineer, Operator, Maintenance engineer Fault diagnosis ANN, SVM, Bayes Classifier
[57], [58], [59] Process signals, Acoustic signals Operator Event prediction ANN
[60], [61], [62] Process signals Operator Soft sensors SVR, ANN, RF
[63], [64], [65] Vibration, Process signals Operator, Maintenance engineer, Scheduler Predictive maintenance RNN, IF
Table 3 shows examples of AI applied to operational activities in the process industry. This table should give an idea of the breadth of use cases, users, relevant data sources, and applicable AI methodologies; however, it is not intended to be a full or systematic examination.
4 Future of Trustworthy (XAI)
Figure 13: Assessing the user’s interaction with XAI [27].
The precise position of each XAI domain and how they relate to the human user are shown in Figure 13. The majority of AI system explanations that are given are usually static and only contain one message [66]. Understanding cannot be attained by explanations alone [67]. Because most existing XAI libraries lack user involvement and customization of explanations, users should be able to explore the system using interactive explanations to gain a better understanding of it. This is a promising research direction for extending the XAI field [67] and [66]. To improve human-machine cooperation and move beyond static explanations, a number of efforts have also been proposed.
Explainable Artificial Intelligence (XAI) has great promise for redefining the relationship between humans and AI systems as it stands at the nexus of technological innovation and societal integration. As AI technologies advance, it is more important than ever to ensure accountability and transparency. Within this framework, XAI becomes a crucial facilitator, entrusted with shedding light on the murky inner workings of AI models and cultivating user confidence. A wide range of breakthroughs are anticipated in XAI, from heightened model transparency and human-centric design principles to regulatory compliance requirements and the rise of hybrid AI systems. XAI approaches will place a high value on user-centric design, providing explanations that are both actionable and understandable. This will foster acceptance and confidence in AI systems.
Moreover, it is anticipated that regulatory frameworks would require the incorporation of XAI in essential applications, guaranteeing compliance with accountability and transparency norms. Future XAI systems will be distinguished by their contextual sensitivity and interactive explanations, which will enable users to interact with AI decisions in real time and adjust to a variety of situations. To guarantee that AI systems follow ethical standards and social values, as well as to democratize access to XAI technologies, efforts must be made to enhance digital literacy and address ethical challenges. Fundamentally, XAI’s success is predicated on its capacity to bridge the communication gap between AI systems and human users, fostering mutual respect, trust, and collaboration in an increasingly AI-dependent world.
This study [68] offers a thorough analysis of Explainable Artificial Intelligence (XAI), focusing on two primary areas of inquiry: general XAI difficulties and research directions, as well as ML life cycle phases-based challenges and research directions. In order to shed light on the significance of formalism, customization of explanations, encouraging reliable AI, interdisciplinary partnerships, interpretability-performance trade-offs, and other topics, the study synthesizes important points from the body of existing literature. The primary contribution is the methodical synthesis and analysis of the body of literature to identify important problems and future directions for XAI research [68]. The research offers a thorough review of the current state of XAI research and provides insightful information for future studies and breakthroughs in the area by structuring the debate around general issues and ML life cycle phases. The primary finding of the study is the identification and clarification of 39 important points that cover a range of issues and potential avenues for future XAI research. The importance of conveying data quality, utilizing human expertise in model development, applying rule extraction for interpretability, addressing security concerns, investigating XAI for reinforcement learning and safety, and taking into account the implications of privacy rights in explanation are just a few of the many topics covered by these points. Furthermore, the paper indicates directions for further research and application by highlighting the potential contributions that XAI may make to a number of fields, including digital forensics, IoT, and 5G.
Figure 14: Issues and Future Research Paths for XAI throughout its Deployment Stage [68].
Deploying machine learning solutions begins the deployment process and continues until we cease utilizing them, possibly even after that. Figure 14 illustrates the XAI research directions and challenges that were explored for this phase.
5 Conclusions
Explainable Artificial Intelligence (XAI) is gaining popularity in a range of fields due to its critical role in addressing critical issues connected to AI adoption. As AI systems become more integrated into society, transparency and interpretability become increasingly important. By offering tools to clarify how AI models make decisions, XAI helps users develop a sense of confidence and comprehension. XAI’s primary objective is to make AI models clear and intelligible. With the help of XAI, the general public will be able to peer inside the black box and comprehend the aspects that affect the AI’s decision-making process. The paper discusses the essential details of XAI and offers a comprehensive overview for a solid understanding. Furthermore, this article discusses in detail the three main application fields of XAI. Lastly, the authors attempt to outline the difficulties in applying XAI and suggest potential future paths.
[…]
Acknowledgements
The authors would like to express their sincere gratitude to everyone who encourages and appreciates their scientific work.
Declarations
Not applicable

### v2

@@
 Abstract
-One of today’s most significant and transformative technologies is the rapidly developing field of artificial intelligence (AI).
-Defined as a computer system that simulates human cognitive processes, AI is present in many aspects of our daily lives, from the self-driving cars on the road to the virtual assistants in our smartphones.
-The term ”black box” is often used to describe artificial intelligence (AI) because some AI systems are so complex and opaque.
-With millions of parameters and layers, these systems—deep neural networks in particular—make it difficult for humans to comprehend how they make judgments.
-Even if AI is capable of producing correct findings, questions of accountability, prejudice, and justice are raised by the opaqueness of its decision-making process.
-AI has a lot of potential, but it also comes with a lot of difficulties and moral dilemmas.
-In the context of explainable artificial intelligence (XAI), trust is crucial as it ensures that AI systems behave consistently, fairly, and ethically.
-In the present article, we explore XAI, reliable XAI, and several practical uses for reliable XAI.
-Once more, we go over the three main components—transparency, explainability, and trustworthiness of XAI—that we determined are pertinent in this situation.
-We present an overview of recent scientific studies that employ trustworthy XAI in relation to these fundamental components, as well as an analysis of how trustworthy XAI is applied in various application fields.
-In the end, trustworthiness is crucial for establishing and maintaining trust between humans and AI systems, facilitating the integration of AI systems into various applications and domains for the benefit of society.
+Artificial Intelligence (AI) is an important part of our everyday lives.
+We use it in self-driving cars and smartphone assistants.
+People often call it a ”black box” because its complex systems, especially deep neural networks, are hard to understand.
+This complexity raises concerns about accountability, bias, and fairness, even though AI can be quite accurate.
+Explainable Artificial Intelligence (XAI) is important for building trust.
+It helps ensure that AI systems work reliably and ethically.
+This article looks at XAI and its three main parts: transparency, explainability, and trustworthiness.
+We will discuss why these components matter in real-life situations.
+We will also review recent studies that show how XAI is used in different fields.
+Ultimately, gaining trust in AI systems is crucial for their successful use in society.
 keywords:
-Artificial Intelligence(AI), Trustworthy XAI, Explainable Artificial Intelligence (XAI), Healthcare, Autonomous Vehicles
+Artificial Intelligence(AI), XAI, Explainable Artificial Intelligence (XAI), Healthcare, Autonomous Vehicles
 1 Introduction
-Philosophers who sought to characterize human thought as the mechanical manipulation of symbols laid the foundations of modern-day artificial intelligence.
-These efforts resulted in the invention of the programmable digital computer [1] in the 1940s.
-Though it is now lost, Alan Turing may have published the first article on the subject of artificial intelligence in 1941, indicating that he was at least thinking about the concept at that time.
-The public was first introduced to Turing’s notion of the Turing test in his seminal work ”Computing Machinery and Intelligence” from 1950 [2].
-In it, Turing cast doubt on the viability of building thinking machines.
-The term artificial intelligence (AI) was first used in 1950 [3] but its general adoption and use in healthcare has been hindered by many shortcomings of the original models.
-The introduction of deep learning in the early 2000s eliminated many of these limitations.
-We are entering a new era of technology where AI can be used in clinical practice through risk assessment models that improve workflow efficiency and diagnostic accuracy.
-AI systems can now analyze complex algorithms and learn on their own.
-The performance of AI systems has improved significantly in recent years.
-These new models expand on their capabilities to include text-image synthesis based on nearly any prompt, whereas previous systems focused primarily on generating facial images.
-Figure 1: Domains throughout which artificial intelligence finds applications. [4]
-The applicability and potential of artificial intelligence (AI) to transform business is already evident in the wide range of areas in which it is applied: In the field of natural language processing (NLP) [5], artificial intelligence (AI) makes it easier for computers to understand and generate human language, enabling tasks such as sentiment analysis, machine translation, and spam filtering.
-Furthermore, computer vision [6] enables computers to comprehend visual information, promoting developments in fields like object identification, facial recognition, and self-driving automobiles.
-Computers can now learn from data thanks to machine learning (ML), which has applications in fraud detection, recommendation systems, predictive analytics, and other fields.
-Robotics [7] is a branch of artificial intelligence that deals with the design, development, and use of machines.
-These machines are used in a variety of industries, including space exploration, manufacturing, healthcare [8], [9], and many more.
-Additionally, the incorporation of AI into business intelligence (BI) [10] signals better data gathering, processing, and display, which promotes data-driven decision-making and increases efficiency.
-AI offers advances in patient outcomes and medical developments in the healthcare industry [11],[12], [13] by assisting in illness diagnosis, therapy development, and tailored care.
-AI’s promise in education is evident in its ability to personalize instruction, engage students, and automate administrative tasks—all of which open the door to more customized learning opportunities.
-Through data-driven tactics, artificial intelligence (AI) in agriculture maximizes crop productivity, lowers expenses, and assures environmental sustainability.
-Similar to this, AI in manufacturing uses task automation and process optimization to increase productivity, efficiency, and quality.
-AI’s influence is felt not just in these domains but also in banking, retail, energy, transportation [14], handwriting detection [15], and government, where it is transforming operations, improving services, and reshaping global industry landscapes.
-Figure 1 illustrates the wide range of industries in which artificial intelligence finds widespread use.
-These industries include retail, security, healthcare, e-commerce, manufacturing, finance, transportation and logistics, and home furnishings.
-These apps depend on reasonably developed AI technologies including ML, NLP, and computer vision.
+The foundations of modern artificial intelligence were laid by philosophers who attempted to define human thought as the mechanical manipulation of symbols, which led to the development of the programmable digital computer [1] in the 1940s.
+Alan Turing may have written the first article on the topic of AI in 1941, though it is now lost, suggesting that he was at least considering the idea at that time.In his groundbreaking essay ”Computing Machinery and Intelligence” from 1950, Turing first presented the idea of the Turing test to the general public [2].
+Turing questioned the feasibility of creating thinking robots in it.
+John McCarthy first used the term artificial intelligence (AI) in 1956 at the Dartmouth Conference[3], but the first models’ numerous flaws have prevented AI from being widely adopted and used in healthcare.
+Many of these limitations were removed with the advent of deep learning in the early 2000s, and we are now entering a new era of technology where AI can be used in clinical practice through risk assessment models that increase diagnostic accuracy and workflow efficiency.
+Performance of AI systems has improved significantly in recent years, and these new models expand on their capabilities to include text-image synthesis based on almost any prompt, whereas previous systems primarily focused on generating facial images.
+Figure 1: Applications of AI across various domains [4]
+The diverse range of fields in which AI is being used already demonstrates its applicability and promise to revolutionize business: AI facilitates activities like sentiment analysis, machine translation, and spam filtering by making it easier for computers to comprehend and produce human language in the discipline of natural language processing (NLP) [5].
+Additionally, computer vision [6] makes it possible for computers to understand visual data, which advances areas like facial recognition, object identification, and self-driving cars.
+Machine learning (ML), which has uses in fraud detection, recommendation systems, predictive analytics, and other domains, has made it possible for computers to learn from data.
+The design, development, and application of machines are the focus of the AI field of robotics [7].
+Many industries, including manufacturing, healthcare, and space exploration, use various machines [8] [9].
+Combining artificial intelligence with business intelligence (BI) [10] improves how businesses collect, process, and visualize data.
+This leads to better decision-making and increased productivity.
+In healthcare, AI helps diagnose diseases, develop treatments, and provide personalized care, which improves patient outcomes [11], [12], [13].
+AI also plays a significant role in education by engaging students, customizing lessons, and automating administrative tasks, resulting in more personalized learning experiences.
+AI in agriculture increases agricultural output, reduces costs, and ensures environmental sustainability through data-driven strategies.
+In a similar vein, AI in manufacturing boosts output, efficiency, and quality through work automation and process optimization.
+AI is changing operations, enhancing services, and changing global industry landscapes in a number of sectors, including banking, retail, energy, transportation [14], handwriting detection [15], and government.
+AI is widely used in a wide number of industries, as seen in Figure 1.
+Retail, security, healthcare, e-commerce, manufacturing, banking, logistics and transportation, and home furnishings are some of these sectors.
+These applications rely on moderately advanced AI technology, such as computer vision, natural language processing, and machine learning.
+The contributions of this research can be stated below:
+  1. 1.
+    Providing an overview of XAI that understands the significance of black box models. For fair and ethical purposes, XAI fosters trust among humans and AI.
+  2. 2.
+    Discussing Deep Learning based systems that will be consistent and trustworthy. Moreover, recent studies from the literature have been reviewed properly.
+  3. 3.
+    Providing guidelines regarding XAI that will be helpful for detecting problems from numerous domains.
+  4. 4.
+    The paper identifies and analyzes three key components of XAI: transparency, explainability, and trustworthiness.
+It details how these elements are essential for understanding and improving AI systems.
 1.1 Third Wave of Artificial Intelligence (3AI)
-The majority of commercial AI technology today is classified as ”narrow AI,” meaning that it consists of extremely specialized systems that excel at a limited number of clearly defined jobs and nothing else.
-Even the most amazing autonomous cars need a combination of limited artificial intelligence algorithms.
-The reliance of modern AI on enormous training data sets is another drawback.
-For example, a three-year-old child can recognize cats from only a few instances, while a standard machine learning algorithm would need to be given tens of thousands of cat photographs before it can recognize them with any degree of precision.
-The idea of ”Third Wave AI” emerges because artificial intelligence has to become more humanlike in a number of ways in order to overcome these constraints and realize its full potential.
+Most current commercial AI technology is called ”narrow AI.” This means these systems are highly specialized and can only perform a few specific tasks.
+For example, even the best self-driving cars rely on limited AI systems.
+Another drawback of today’s AI is its reliance on large training data sets.
+A typical machine learning program needs tens of thousands of cat photos to recognize cats accurately, while a three-year-old child can do this with just a few examples.
+The idea of ”Third Wave AI” comes from the need for AI to become more humanlike in various ways to overcome these limitations and achieve its full potential.
 Figure 2: Past, Present, and Future of AI waves. [16]
-According to the Defense Advanced Research Projects Agency (DARPA) [17], third-wave AI systems will be able to understand context, apply that contextual awareness with common sense, and adapt to changing conditions.
-This will make it possible for AI systems and human users to connect in a more organic and intuitive way [17].
-One of the few DARPA projects underway, XAI is anticipated to pave the way for ”third-wave AI systems,” in which computers comprehend the context and surroundings in which they function and gradually develop the underlying explanatory models necessary to describe occurrences in the actual world.
-  • First Wave AI was centered on rules, logic, and constructed knowledge.
-  • Big data, statistical learning, and probabilistic techniques were introduced by Second Wave AI.
-  • The goal of third-wave AI is to develop common sense and contextual adaption skills.
-According to Tractica [16], the worldwide market for artificial intelligence software is predicted to generate revenues of 118.6 billion by 2025, up from around 9.5 billion US dollars in 2018.
-The focus of this data is on developing AI systems that not only perform tasks accurately but also provide explanations of their decisions and actions in ways that humans can understand.
-The third wave of AI refers to the evolution of AI technologies beyond traditional machine learning methods towards more advanced approaches that integrate reasoning, context awareness, and human-like understanding.
-This wave aims to develop AI systems capable of understanding and interacting with the world in more nuanced and sophisticated ways, often drawing inspiration from cognitive science and neuroscience.
-Explainable AI (XAI), on the other hand, focuses on improving the transparency and interpretability of AI systems, especially machine learning models, so that humans can understand and trust the decisions made by these systems.
-XAI techniques aim to provide explanations of AI predictions and actions, enabling users, including developers, regulators, and end users, to understand the underlying considerations and factors that influence AI outcomes.
-The relationship between the third wave of AI and XAI lies in their shared goal of creating AI systems that are not only capable of making accurate predictions or decisions but also capable of providing meaningful explanations for those predictions or decisions.
-By incorporating XAI techniques into the development of third-wave AI systems, developers can ensure that these systems are not only advanced in their capabilities but also transparent and interpretable, fostering trust and acceptance among users.
+According to the Defense Advanced Research Projects Agency (DARPA) [17], third-wave AI systems will understand the context of situations, use common sense, and adapt to changes.
+This will create more natural and intuitive connections between AI systems and people [17].
+One of DARPA’s active projects, called XAI, aims to develop these third-wave AI systems.
+These computers will learn about their environments and the contexts in which they work.
+They will also build explanations needed to clarify real-world events.
+  • First Wave AI focused on rules, logic, and built knowledge.
+  • Second Wave AI introduced big data, statistical learning, and probabilistic techniques.
+  • The goal of third-wave AI is to develop common sense and the ability to adapt to different contexts.
+Tractica [16] predicts that the global market for AI software will grow from about 9.5 billion US dollars in 2018 to 118.6 billion by 2025.
+This data aims to develop AI systems that can perform tasks accurately while providing explanations that people can understand.
+The term ”third wave” refers to the advancement of AI technologies beyond traditional machine learning.
+This new phase focuses on creating more advanced systems that can understand context, reason, and think similarly to humans.
+It draws inspiration from cognitive science and neuroscience, aiming to build AI that can engage with the world in more complex and detailed ways.
+XAI, or Explainable Artificial Intelligence, focuses on making AI systems, especially machine learning models, easier to understand.
+The goal is to help people trust the decisions these systems make.
+XAI techniques work to explain how AI makes predictions and why it behaves the way it does.
+This helps users, such as developers, regulators, and everyday users, grasp the key factors that affect AI results.
+Developing AI systems that can make accurate predictions or decisions and explain why they did so is a key goal of both the third wave of AI and explainable AI.
+By using XAI methods in their design, developers can ensure that these advanced AI systems are not only effective but also easy to understand.
+This will help build user trust and acceptance.
 1.2 Concept of Explainable AI
-Artificial Intelligence is often criticized for being hard to explain.
-Many opponents argue that it is hard to trust the results of an AI if one does not know how it can arrive at a certain result or conclusion.
-This issue becomes particularly problematic when AI-based programs and systems are unable to accomplish their goals.
-Developing explainability is necessary to boost public confidence in the computational execution.
-Figure 3 shows the AI in use today and the AI expected to be used tomorrow.
-If we are to hold the system accountable, we must take steps to address and minimize input inefficiencies [18].
-Artificial intelligence is known to suffer from the ”black box” syndrome due to a lack of understanding of how the system works.
-This has ramifications for human confidentiality, arbitrary discrimination, obfuscation, and legitimacy.
-There are often underlying prejudices and tendencies present along with this lack of openness.
-By enhancing users’ comprehension of how AI-powered systems reason, XAI seeks to improve their performance.
-The goal of Transparent AI is to make artificial intelligence (AI) safer and more accessible than ever before [18].
-Therefore, each of the TAI’s diverse features has to be considered separately, and their many facets need to be talked about.
-AI’s use of machine learning (ML) techniques falls into two categories: white box and black box [19].
-The results of a white box model are understandable to a subject matter expert.
-In contrast, a black box model is very difficult to explain and can be confusing even to domain experts [20].
-The three criteria of interpretability, explainability, and transparency are adhered to by the XAI algorithm [21].
-A model is transparent when the underlying principles of the machine learning model and decision-making can be understood ”when the process of extracting model parameters from training data and generating labels from test data is described and justified by the designers of the approach.
-” Because communicating in a way that people can understand is called interpretability [22].
-Currently, there is no universally accepted definition of the concept of explainability, but its importance is recognized.
-An alternative is ”a collection of interpretable features of a domain that can help generate a decision (such as classification or regression) about a given example.” If an algorithm follows these guidelines, it provides a basis for recording and validating the decision, improving the algorithm, and discovering new information.
+AI often struggles with what is called the ”black box” problem because users do not understand how it works.
+This can lead to issues like lack of trust, confusion, unfair treatment, and violations of privacy.
+AI systems can also have hidden biases.
+Explainable AI (XAI) aims to make AI systems easier to understand, helping users know how they make decisions.
+The goal of XAI is to make AI safer and more user-friendly.
+Therefore, we need to look at each part of AI individually and discuss its different aspects [18].
+There are two main types of machine learning (ML) techniques: white box and black box models [19].
+Experts can easily understand the results of a white box model.
+However, even specialists may find it hard to grasp the results of a black box model [20].
+The XAI algorithm [21] follows three key principles: interpretability, explainability, and transparency.
+A model is transparent when it clearly explains how it gets its results from training data and how it generates labels from test data.
+Interpretability means being able to explain findings in a way that others understand [22].
+While there’s no single accepted definition of explainability, its value is recognized.
+One definition describes it as a set of clear features that help make decisions, like classifying or predicting outcomes for specific cases.
+An algorithm that meets these standards helps document and verify decisions and improves itself based on new data.
 Figure 3: Explainable Artificial Intelligence (XAI): A look at AI now and tomorrow. [18]
-From an explainability perspective, academics are actively studying intelligent systems, which is a key topic.
-Formal compliance may occasionally depend on the system’s understanding.
-Many black box algorithms, as shown in Figure 4, present a trade-off between incredible learning performance (accuracy) and provable AI.
-Figure 4: Performance of artificial intelligence vs explainability. [23]
+Researchers are studying intelligent systems to understand them better.
+This is an important topic.
+Sometimes, a system needs to understand its own workings to comply with rules.
+Many complex algorithms, shown in Figure 4, balance achieving high accuracy with being explainable.
+Figure 4: Trade-off between AI model accuracy and explainability, highlighting the challenge of balancing performance with interpretability. [23]
 1.3 Classification Tree of XAI
@@
 Transparent approaches can be useful when the internal feature correlations are highly complex or linear.
-A comprehensive classification of different XAI methods and approaches related to different types of data is shown in Figure LABEL:Classification [lertvittaakumjorn2019human].
-Figure 5: XAI categorization according on data type [24]
+A comprehensive classification of different XAI methods and approaches related to different types of data is shown in Figure 5.
+Figure 5: Categorization of Explainable AI (XAI) techniques based on data type, illustrating differences between transparent and posthoc approaches [24]
 Posterior approaches are useful for interpreting the complexity of a model, especially when there are nonlinear relationships or high data complexity.
@@
 1.4 Definition of Transparency in Artificial Intelligence
-In explainable artificial intelligence (XAI), transparency refers to the ability of an AI system to provide understandable justification for its decisions and actions [25].
-In fact, transparency is one of the key components of explainable artificial intelligence (XAI).
-In many real-world applications, particularly those with large social implications, deciphering the reasoning behind an AI system’s choice is just as crucial as the decision itself.
-Assuring transparency in XAI helps prevent AI systems from being viewed as ”black boxes” and instead as instruments that enable users to get insightful and informative responses [26].
-The general transparency of AI systems is further enhanced by the insights offered by XAI approaches.
-Users have the capacity to scrutinize the decision-making procedures, detect any partialities, and evaluate the dependability and equity of the model’s results [27].
-Transparent systems are essential for guaranteeing accountability and ethical concerns in fields where the implications of AI choices might have large repercussions, such as healthcare, finance, and autonomous cars.
-XAI approaches enable users to find patterns, comprehend relationships, and discover any flaws or biases by providing insightful information about the inner workings of AI models [27].
-Stakeholders are better equipped to make judgments, confirm that the model’s predictions are accurate, and take necessary action as a result of the improved openness.
-The examination of the ethical criteria revealed a correlation between explainability and transparency and a number of other quality needs.
-The nine quality standards pertaining to explainability and openness are shown in Figure 6.
-Figure 6: Qualitative standards of explainability and openness + Assistance; – Disagreements. [28]
+Transparency in XAI is the capacity of an AI system to make its decisions and actions understandable through explanations [25].
+Transparency is one of the most important aspects of XAI, rendering AI decisions interpretable and justified.
+Transparency allows decision-making processes to be closely examined by stakeholders, mitigating risks in applications with societal impact, such as healthcare and finance [26].
+The explainability provided by XAI techniques also increases the overall transparency of AI systems.
+The users are able to examine the decision-making procedure, identify any bias, and analyze the reliability and fairness of the output of the model [27].
+Transparent solutions are necessary in areas like the medical field, banking, and autonomous vehicles, where AI-decisions can have significant effects.
+By providing meaningful information about the internal processes of AI models, XAI methods assist users in recognizing patterns, comprehending relationships, and revealing biases or errors [27].
+Due to the heightened transparency, stakeholders can more effectively develop opinions, ensure the predictions made by the model are accurate, and take the necessary action.
+The examination of ethical criteria showed a link between explainability, transparency, and other quality needs.
+Figure 6 displays nine quality standards related to explainability and openness.
+The key standards for AI transparency are marked with “O.” For example, O2 focuses on how to interpret models, O15 highlights the importance of traceability, and O5 and O12 ensure that users understand the information.
+By following these standards, AI models can provide clear and convincing explanations for their outcomes.
+Keeping these standards improves accountability and helps reduce bias in AI systems.
+Figure 6: Key qualitative standards (O1-O15) related to explainability and transparency in AI systems, addressing user comprehension, interpretability, and traceability. [28]
 The growth of AI systems’ explainability and transparency is facilitated by their understandability.
@@
 1.5 Transparency Vs Explainability in AI
-Transparency and explainability can be compared [29].
-When an AI is transparent, the ”basic elements of data and decisions must be available for inspection during and after AI use,” according to McLarney et al.
+Explainability and transparency are similar concepts [29].
+According to McLarney et al.
+[30], a transparent AI necessitates that ”Basic elements of data and decisions must be available for inspection during and after AI use.” Transparency is achieved when users have access to their data or can understand how decisions are made.
+On the other hand, explainability seeks to reveal the reasons for AI’s successes or failures and demonstrate how it utilizes the knowledge and judgment of those it affects.
+It provides a rational justification for the actions of the AI.
+Users must clearly know what data is collected, how the AI interprets this data, and how it produces reliable outcomes for each affected individual.
+This straightforward explanation overlooks the challenges we face when trying to clarify ”black box” algorithms, the context that may be omitted, and the accuracy needed when offering understandable explanations to customers.
+Therefore, the question arises: is having minimal explainability preferable to having none at all?
 [30].
-Transparency exists when a user can observe how decisions are made or have access to their data.
-Explainability, on the other hand, is about understanding why AI succeeds or fails and revealing how it draws on the knowledge and decision-making processes of the people it will affect.
-It provides a rational explanation for the AI’s actions.
-Users need to be able to understand what data is being collected, how the AI program processes it, and how it produces trustworthy outcomes for each individual affected.
-This straightforward explanation ignores the challenges we confront in simplifying ”black box” algorithms, the context that is lost, and the accuracy needed when giving consumers clear explanations.
-The question thus becomes, is minimal explainability preferable to nothing?
-[30].
-Other important factors to consider include the belief that explanations can adequately account for the dynamic nature of the rich information ecosystem and the appropriateness of dealing with anomalies.
+Additionally, the belief that explanations can adequately address the dynamic nature of the rich information ecosystem and the appropriateness of managing anomalies are also vital factors to consider.
 Figure 7: Output from the Bing search engine’s conversation feature explaining a failure. a partial screenshot taken using an Android smartphone on March 2, 2023. [17]
-Interestingly, while certain AI algorithms analyze data automatically, an increasing number of AI systems are designed to explain how their algorithms work and the reasoning behind certain decisions [17].
-For example, the Bing search engine’s conversation mode offers brief explanations of how it works (Fig.
+It’s interesting to note that although certain AI algorithms evaluate data automatically, more and more AI systems are made to explain how their algorithms operate and the logic behind specific choices [17].
+For instance, the conversation mode of the Bing search engine provides succinct explanations of its operation (Fig.
 7).
-End users may occasionally find these explanations enough, but occasionally they may get confused about how an AI arrived at a specific conclusion or behaved in a specific way.
-It is impractical to expect people to become more computer literate when they are more perplexed by the explanation that is provided [17].
-Rather, we need to either enhance the AI system or the justification.
+Sometimes, end users might find these explanations sufficient, but other times, they would be perplexed as to how an AI came to a particular conclusion or acted in a particular manner.
+When individuals are more confused by the explanation given, it is unrealistic to expect them to become more computer-literate [17].
+Instead, we must improve the justification of the AI system.
 1.6 Definition of Trustworthiness in Artificial Intelligence
-Artificial intelligence (AI) systems that incorporate trustworthiness must take a multifaceted approach that takes organizational, ethical, and technical factors into account.
-Establishing criteria for assessing trustworthiness is the first step in this process.
-These measures should include accountability, security, privacy, openness, fairness, and ethical compliance.
-Foundational elements include transparent algorithms that provide intelligible justifications for AI-driven judgments and high-quality, impartial data.
-Strong security measures and privacy-preserving strategies protect sensitive data and fight off online attacks.
-Responsible AI usage is encouraged through the establishment of accountability systems and adherence to moral principles and governance structures.
-User-centric design, ongoing observation, and training guarantee that AI systems satisfy users’ demands while developing over time to retain their credibility.
-Organizations may create trustworthy, transparent, equitable, and ethical systems that inspire confidence in both users and stakeholders by incorporating these principles into all phases of the AI lifecycle.
-Figure 8: The three primary elements of a reliable AI. [31]
-The three elements depicted in Figure 8 - algorithmic ethics, data ethics, and practice ethics - come together to form trustworthy AI.
-These elements offer an abstraction level for ethical issues that is data-centric [31].
-Numerous unresolved problems arise when trying to address ethical concerns with AI systems.
-In the research paper [31], authors describe the requirements for Trustworthy AI where the mentioned:
+Creating trustworthy AI systems requires a careful strategy that looks at organizational, ethical, and technical factors.
+The first step is to set clear standards for trustworthiness.
+These standards should include accountability, security, privacy, transparency, fairness, and ethical behavior.
+Using high-quality, unbiased data and clear algorithms that explain AI decisions is essential.
+Strong security measures and privacy practices protect sensitive information from cyberattacks.
+It’s important to create accountability frameworks and follow ethical guidelines to ensure responsible AI use.
+By focusing on user needs and constantly monitoring and updating the systems, AI can stay reliable over time.
+Applying these principles across all stages of the AI process allows organizations to develop systems that are explainable, equitable, ethical, and robust, which fosters stakeholder and user trust.
+Figure 8: The three key components of XAI: Algorithmic Ethics, Data Ethics, and Practice Ethics. [31]
+The three elements illustrated in Figure 8—algorithmic ethics, data ethics, and practice ethics—intersect to create responsible AI.
+These elements define a data-centered way to handle ethical issues [31].
+However, several open challenges still remain with respect to dealing with ethical issues in AI systems.
+In the work [31], authors give the vision of Trustworthy AI, which mentions that:
   1. 1)
-    Human agency and oversight: AI systems must uphold the values of human freedom.
-AI systems must promote user agency, uphold fundamental rights, and enable human oversight to realize a democratic and equal society.
+    Human agency and oversight: AI systems have to enable human freedom.
+They need to facilitate user choice, safeguard fundamental rights, and enable human control.
+This will assist in developing an equitable and just society.
   2. 2)
-    Security and technical robustness: Security and technical robustness lead to the prevention of harm.
-For an AI system to function reliably while minimizing harm, risks must be taken into account during development.
-This must cover any changes in the working environment as well as possible attacks on the system by adversaries.
+    Security and technical robustness: Security and technical proficiency are crucial to prevent damage.
+In order to enable an AI system to operate efficiently and reduce risks, its creators ought to consider potential risks when designing it.
+They range from environmental alterations where the system will operate to attacks by malicious individuals.
   3. 3)
-    Data protection and data governance: As a fundamental right that is particularly affected by the extensive data collection required by AI systems, data protection is also closely related to preventing harm.
-Preventing privacy harm also requires data governance that addresses the quality and integrity of the data used, its relevance, access protocols, and the ability to process it in a privacy-preserving way.
+    Data protection and data governance: Privacy is a fundamental right that has been highly compromised with the vast amounts of data that artificial intelligence systems gather.
+It is necessary to protect individual privacy to prevent potential harm.
+For this purpose, robust data governance must be in place.
+This includes making sure that the information being utilized is precise and applicable.
+Furthermore, there is a necessity to establish definite rules for data access and how data must be treated while maintaining the integrity of privacy.
   4. 4)
-    Transparency: Explainability and transparency are closely related requirements.
-The aim is to make all relevant aspects of an AI system transparent, including data, technology, and business models.
-In the era of ubiquitous computing, transparency is essential to support large-scale data collection and its benefits to consumers.
+    Transparency: Explainability and transparency are pretty much dependent on each other.
+The key objective is to make data, technology, and business models clear.
+In today’s age, which is the age of pervasive technology, transparency has become a must.
+It aids customers in comprehending the huge volumes of data collected and the ensuing benefits.
   5. 5)
-    Diversity, non-discrimination, and fairness: Achieving trustworthy AI requires enabling inclusivity and diversity throughout AI systems.
-This is important not only to consider and involve all affected parties but also to ensure fair access and treatment.
-Fairness and this need go hand in hand.
+    Fairness, diversity, and nondiscrimination: Including several voices in AI systems is vital to achieve XAI.
+We need to involve all individuals who may be affected to ensure equal treatment and access.
+Fairness and this requirement come hand in hand.
   6. 6)
-    Social and environmental welfare: In the spirit of justice and harm prevention, the environment and the broader community should be considered as stakeholders.
-Research into AI solutions to address global challenges should be promoted, making AI systems more environmentally friendly and sustainable.
-AI systems need to benefit everyone, including those who will come after us.
+    Social and environmental welfare: We have to think about the environment and community in seeking justice and doing no harm.
+We should finance research on AI solutions to global issues.
+This will make AI systems environmentally friendly and sustainable.
+AI is supposed to be for the good of all people, including future generations..
   7. 7)
-    Accountability: The idea of fairness and the need for accountability go hand in hand: we need mechanisms to ensure accountability and accountability for AI systems and the results they produce, both during and after their development, application, and deployment.
+    Accountability: Accountability and fairness are essential in the context of AI.
+We need to have systems of holding AI systems accountable for their actions and generated results.
+Accountability needs to be a constituent part of AI development, deployment, and use, both during and following the activities.
+1.7 Impact of XAI on Zero Trust Architecture (ZTA)
+Zero Trust Architecture (ZTA) is a security system that always checks every request, no matter where it comes from, and does not assume trust automatically.
+Explainable AI (XAI) helps ZTA by ensuring that decisions made by AI in security are clear and reasonable.
+XAI is especially useful in identity verification, access control, and spotting unusual behavior.
+AI models analyze how users behave and identify any suspicious activities.
+By adding explainability, security analysts can better understand and confirm AI-driven security rules.
+This reduces false alarms and speeds up response times.
+For example, AI-driven network monitoring systems that use ZTA principles can explain why a specific access attempt looks suspicious.
+This explanation builds trust in automated cybersecurity decisions [32].
 1.8 An Overview of Necessities for Reliable AI
-The conditions for reliable AI are still unclear and are addressed incongruously by many institutions and groups, despite contentious social debates over it.
-The principles of Fairness, Responsibility, Accuracy, Verifiability, and Accountability in Machine Learning (FAT-ML) include accountability, explainability, verifiability, and fairness at a global level [32].
-Among the numerous needs under review, explainability, fairness, privacy, and robustness will all be covered in this study (Table 1).
+Despite heated societal discussions, the requirements for trustworthy AI remain ambiguous and are handled inconsistently by numerous organizations and organizations.
+Globally, accountability, explainability, verifiability, and fairness are all part of the Fairness, Responsibility, Accuracy, Verifiability, and Accountability in Machine Learning (FAT-ML) principles [33].
+Explainability, fairness, privacy, and robustness are just a few of the many needs that will be examined in this study (Table 1).
 Table 1: Conditions necessary for trustworthy artificial intelligence (AI)
@@
 Robustness The AI model can fend against outside threats while continuing to operate correctly.
-2 Trustworthy XAI Vs AI
-A paradigm change in the field of artificial intelligence (AI) has been brought about with the introduction of Trustworthy Explainable AI (XAI).
-Oftentimes, conventional AI systems operate as opaque black boxes, making it challenging for users to comprehend the decision-making process.
-On the other hand, Trustworthy XAI tackles significant concerns surrounding the adoption of AI by focusing on accountability, interpretability, and transparency.
-Reputable XAI seeks to build consumers’ confidence by providing clear justifications for its choices.
-Users are able to evaluate the fairness and dependability of AI-driven results because to this transparency.
-While typical AI systems are capable of producing precise forecasts or suggestions, they do not have the openness required to establish credibility.
-The way that reliable XAI and traditional AI make decisions is what sets them apart from one another.
-Although AI systems have the potential to produce precise forecasts or suggestions, Reputable XAI highlights the necessity of providing justifications that clarify the processes that lead to these results.
-Users may evaluate the fairness and dependability of AI-driven outcomes thanks to the clear justifications that trustworthy XAI systems give for their decisions.
-Furthermore, trustworthy XAI takes into account more ethical factors than just explanations.
-Incorporating the Fairness, Accountability, and Transparency (FAT) principles into AI development procedures, guarantees that AI systems abide by moral and legal requirements.
-Trustworthy XAI seeks to reduce biases, discrimination, and other possible negative effects of AI technology by placing a high priority on ethical norms.
-An approach to AI development known as trustworthy AI places a high value on user safety and openness.
-Since no model is flawless, trustworthy AI developers take care to explain to clients and the wider public how the technology was developed, its intended applications, and its limitations.
+2 XAI Vs AI
+XAI improves AI systems by focusing on transparency, clear explanations, and accountability.
+It offers understandable reasons for decisions, which helps users trust the system and makes it easier to assess fairness compared to traditional ”black box” methods.
+The main difference between reliable XAI and traditional AI is how they make decisions.
+While AI can give accurate forecasts or suggestions, reliable XAI emphasizes the need to explain the steps that lead to these results.
+Clear explanations from XAI systems allow users to judge the fairness and reliability of AI-generated outcomes.
+To improve security and maintain transparency in AI-driven cybersecurity, we need to integrate XAI into Zero Trust Architecture (ZTA).
+When explainability methods clarify why certain decisions are made, people can better understand and trust the AI-driven access control and behavioral analytics in ZTA.
+As we face compliance and operational challenges, future cybersecurity frameworks will rely more on AI automation.
+It will be essential to ensure that these AI systems can be easily explained [34].
+XAI focuses on more than just providing explanations; it also considers ethical issues.
+AI development processes that follow the principles of Fairness, Accountability, and Transparency (FAT) help ensure that AI systems meet ethical and legal standards.
+By prioritizing ethical standards, XAI aims to reduce biases, discrimination, and other harmful effects of AI technology.
+Trustworthy AI is an approach that emphasizes user safety and transparency.
+Responsible AI developers clearly explain to clients and the public how the technology works, what it is meant for, and its limitations, since no model is perfect.
 Table 2: Seven Requirements to Meet in Order to Develop Reliable AI
@@
 Data Governance and Privacy Individuals should be in total control of the information that is about them, and information about them should not be used against them Notification and information access rights regarding the logic used in automated processes Art 13, 14, and 15
-Transparency Systems using artificial intelligence ought to be transparent and traceable The right to get clarification Recital 71
+Transparency Systems using AI ought to be transparent and traceable The right to get clarification Recital 71
 Diversity and Fairness AI systems have to provide accessibility and take into account the whole spectrum of human capacities, requirements, and standards Right to not have decisions made only by machines Art 22
@@
 Accountability Establishing procedures to guarantee that AI systems and their outcomes are held accountable is essential Right to be informed when decisions are made only by machines Art 13, 14
-3 Applications of Trustworthy XAI
-Authentic Explainable Artificial Intelligence (XAI) has numerous uses in sectors where accountability, interpretability, and transparency are essential.
+3 Applications of XAI
+Authentic XAI has numerous uses in sectors where accountability, interpretability, and transparency are essential.
 XAI can provide an explanation for a diagnosis or therapy recommendation in medical diagnosis and recommendation systems.
@@
 We shall concentrate on a few particular applications in this section and go into detail about them.
-3.1 Application of Trustworthy XAI in Medical Science
-The field of artificial intelligence (AI) is rapidly growing on a global scale.
-The potential uses of artificial intelligence in healthcare are a hot topic for research [33].
-There are many opportunities to use AI technology in the healthcare sector, where people’s lives and well-being are in danger, because of its essential relevance and the enormous quantity of digital medical data that have been gathered [34].
-Artificial intelligence (AI) has made it possible to accomplish tasks quickly that were before unfeasible for traditional technologies.
-Trustworthy AI is a huge concern these days.
-Since incidents involving AI-powered chatbots such as Tay and Iruda [35], there has been an increase in interest in the topic of whether an AI’s judgment and decision-making system is reliable.
-The credibility of AI in the medical and healthcare areas requires more investigation.
-Clinical decision support systems (CDSS) in the medical field use AI technology to aid with important medical duties such as diagnosis and therapy planning [36].
-Misuse can have serious consequences in areas where lives are at stake, even if the scope of use is limited to assisting healthcare practitioners.
-False alarms, for example, which occur often in scenarios involving urgent patients, may exhaust medical personnel.
-The study [37] adds significantly to the field of medical skin lesion diagnostics in a number of ways.
-Before anything else, it adapts an existing eXplainable Artificial Intelligence (XAI) technique to increase user confidence and trust in AI decision-making systems.
-This modification entails describing an AI model that is skilled at differentiating between different kinds of skin lesions.
-Synthetic exemplar and counter-exemplar images are used to create explanations that illustrate the important characteristics that influence classification choices.
-This research [37] is based on training a deep learning classifier with the ISIC 2019 dataset using the ResNet architecture.
-This enables practitioners to use the explanations offered to reason effectively.
-All things considered, the study’s original contributions are found in the way it refined and assessed the XAI technique in an actual medical setting, examined the latent space and carried out an extensive user study to gauge the efficacy of the explanations, especially among subject matter experts.
-This research paper [38] addresses the challenge of recognizing brain tumors in MRI images by merging two powerful algorithms: fuzzy C-means (the FCM method) and Artificial Neural Network (ANN).
-The authors want to increase the segmentation process’s interpretability as well as the accuracy of tumor identification by merging these techniques.
-Their main objective is to improve medical decision-support system tools so that physicians can diagnose brain tumors more accurately.
-This discovery has two main benefits: first, it improves the ability to identify brain cancers in medical imaging with more accuracy, which is important for early diagnosis and treatment.
-Second, the researchers make the decisions made by their models more transparent and intelligible to patients and medical experts by integrating explainable AI principles into the segmentation process.
-In the end, this improved interpretability raises the general level of reliability and acceptance of AI-driven systems for medical picture segmentation in clinical settings.
-In another research [39], the field of computational pathology—which uses machine learning and artificial intelligence (AI) to diagnose whole slide images (WSIs)—is discussed.
-Because artificial intelligence is opaque, there are doubts regarding its reliability despite the technology’s potential to improve efficiency and accuracy.
-The paper suggests employing explainable AI (xAI) techniques, which can shed light on the choices made by AI algorithms, to allay these worries.
-Computational pathology systems become more clear and dependable with the addition of xAI, especially when it comes to crucial activities like pathology diagnosis.
-Additionally, it presents HistoMapr-Breast, a software program with xAI capabilities intended for breast core biopsies.
-In a recent research [40], it is discussed how crucial it is to guarantee the correctness and resilience of AI-based systems in the healthcare industry, especially with regard to their interpretability and defense against hostile attacks.
-As AI systems are used in medical contexts more frequently, it is important to confirm that the predictions they are generating are based on accurate features.
-Numerous model interpretability and explainability techniques have been put out in an effort to overcome this.
-This work shows that even with strong training, adversarial attacks can affect the explainability of a model.
-The authors also introduce two attack classifiers: one to determine the type of attack and another to differentiate between benign and opposed inputs.
-This research paper [41] studies the rising area of explainable machine learning in cardiology.
-It addresses challenges with the interpretability of complicated prediction models and their ramifications for key healthcare choices.
-This study delves into the underlying principles and tactics of explainable machine learning, providing cardiologists with a greater grasp of the approach’s strengths and limitations.
-The research intends to help decision-making processes in clinical settings by providing a rule of thumb for deciding between interpretable and black-box models.
-This will ultimately improve patient outcomes while maintaining accountability and transparency in model predictions.
-Figure 9 was created by training a single new decision tree based on the random forest model predictions.
-The global tree diagram depicts the overall operation of the random forest.
-Individual predictions can then be investigated by tracking the patient’s progress through the global tree.
-This type of explanation has the advantage of making it easy to understand both the overall operation of the model and the reasoning behind each particular forecast.
-Decision trees are ideal for disciplines such as cardiology because they provide rule-based reasoning similar to clinical decision criteria.
-Figure 9: A random forest-based model for heart disease prediction is described, using both local and global decision trees.
-Because decision trees are evaluated from top to bottom, the global diagram shows that the model begins by deciding if patients’ thallium stress test results are normal.
-The model investigates the patient’s ST depression if the thallium stress test discloses an issue, and so on.
-The local graphic depicts the path a patient took to fall from the tree, showing the reasoning for each individual prognosis.
-The patient was less than 54.5 years old, had a maximum heart rate of higher than, and had a normal thallium stress test result, indicating a very low chance of cardiac disease [41].
-Figure 10 depicts the LIME explanations for our heart failure model’s two local predictions.
-The authors demonstrate how these predictions are integrated as a clinical decision support tool in Epic, an electronic health record created with doctors in mind (Epic Systems Corporation, Verona, Wisconsin, USA).37 This type of explainability approach has the advantage of providing detailed explanations for the clinical factors that impact each prediction.
-It is noteworthy that this kind of explanation can be incorporated into an EHR, as this could enhance a black-box model’s actionability by allowing forecasts and clear explanations to be effortlessly integrated into clinical workflow.
+3.1 Application of XAI in Medical Science
+The field of artificial intelligence (AI) is rapidly growing on a global scale, particularly in healthcare, which is a hot topic for research [35].
+There are numerous opportunities to utilize AI technology in the healthcare sector, where the well-being of individuals is at stake, due to its significant relevance and the vast amounts of digital medical data that have been collected [36].
+AI has enabled us to perform tasks quickly that were previously unfeasible with traditional technologies.
+The trustworthiness and openness of AI systems are becoming increasingly important, especially in areas like healthcare.
+As AI is used more in medical decision-making, people are worried about how reliable and understandable its results are.
+These worries highlight the need to evaluate AI models carefully to make sure their predictions are based on important and verifiable factors.
+In critical situations like medicine, proving that AI systems are credible is vital for their safe and effective use.
+In the medical field, clinical decision support systems (CDSS) utilize AI technology to assist healthcare professionals with critical tasks such as diagnosis and treatment planning [37].
+While these systems aim to support healthcare practitioners, misuse can have severe consequences in situations where lives are at risk.
+For example, false alarms, which are common in scenarios involving urgent patients, can lead to exhaustion among medical personnel.
+The study [38] significantly contributes to medical skin lesion diagnostics in several ways.
+First, it modifies an existing explainable AI (XAI) technique to boost user confidence and trust in AI systems.
+This change involves developing an AI model that can distinguish between different types of skin lesions.
+The study uses synthetic examples and counter-examples to create explanations that highlight the key features influencing classification decisions.
+The research [38] trains a deep learning classifier with the ISIC 2019 dataset using the ResNet architecture.
+This allows professionals to use the explanations to reason effectively.
+Overall, the study’s main contributions lie in its refinement and evaluation of the XAI technique in a real-world medical setting, its analysis of the latent space, and its thorough user study to assess how effective the explanations are, particularly among experts in the field.
+This research paper [39] discusses how to recognize brain tumors in MRI images using two effective algorithms: fuzzy C-means (FCM) and Artificial Neural Network (ANN).
+The authors aim to make the tumor segmentation process more understandable and improve accuracy in identifying tumors.
+Their main goal is to enhance tools that help doctors diagnose brain tumors more accurately.
+This research offers two key benefits.
+First, it helps identify brain cancers in medical images more precisely, which is crucial for early diagnosis and treatment.
+Second, by incorporating XAI principles into the segmentation process, the researchers make their models’ decisions clearer and easier to understand for patients and medical experts.
+In summary, this increased clarity boosts the overall trust and acceptance of AI-driven systems in medical image analysis within clinical settings.
+This study [40] discusses how AI and machine learning can help diagnose whole slide images (WSIs) in pathology.
+While AI can improve accuracy and efficiency, concerns exist about its reliability because it can be hard to understand.
+To address these issues, the article suggests using explainable AI methods, which help clarify how AI makes decisions.
+By adding XAI, pathology systems become more transparent and trustworthy, especially for critical tasks like diagnosing diseases.
+The study also introduces HistoMapr-Breast, a software tool that uses XAI to assist with breast core biopsies.
+A recent study examines the importance of making sure AI systems in healthcare are accurate and strong, especially regarding how easy they are to understand and how well they can resist attacks [41].
+As AI becomes more common in medical settings, it’s crucial to verify that the predictions these systems make rely on trustworthy features.
+To tackle this challenge, researchers have proposed various methods to improve model interpretability and explainability.
+The study shows that adversarial attacks can affect a model’s explainability, even when the model has strong training.
+Additionally, the authors introduce two types of attack classifiers: one that tells apart harmless and harmful inputs, and another that determines the nature of the attack.
+This research paper [42] looks at explainable machine learning in cardiology.
+It discusses the challenges of understanding complex prediction models and how these models affect important healthcare decisions.
+The study explains the main ideas and methods of explainable machine learning, helping cardiologists understand the benefits and limitations of this approach.
+The goal is to improve decision-making in clinical settings by offering guidance on when to use easy-to-understand models versus complex ones.
+This can help improve patient outcomes while ensuring accountability and transparency in predictions.
+Figure 9 shows a decision tree created from the predictions of a random forest model.
+This global tree diagram illustrates how the random forest works overall.
+By following a patient’s path through the tree, individual predictions can be examined.
+This type of explanation is beneficial because it clarifies both the general functioning of the model and the reasoning behind specific predictions.
+Decision trees are suitable for fields like cardiology because they use rule-based reasoning similar to clinical decision guidelines.
+Figure 9: A model using random forests predicts heart disease by analyzing both local and global decision trees.
+The global diagram starts by examining whether a patient’s thallium stress test results are normal.
+If the test shows a problem, the model looks at the patient’s ST depression next.
+The local graphic shows the specific pathway a patient took in the model, explaining the reasons for their individual prediction.
+For example, a patient under 54.5 years old, with a maximum heart rate that is high and normal thallium stress test results, has a very low chance of having heart disease [42].
+Figure 10 shows the LIME explanations for our heart failure model’s two local predictions.
+The authors explain how these predictions serve as a clinical decision support tool in Epic, which is an electronic health record designed for doctors (Epic Systems Corporation, Verona, Wisconsin, USA).
+This kind of explanation helps to clarify the clinical factors that affect each prediction.
+Importantly, this type of explanation can be added to an EHR, which may improve the practical use of a complex model by making forecasts and clear explanations easy to integrate into clinical work.
 Figure 10: Heart disease prediction explanation produced using Local Interpretable Model-Agnostic Explanations (LIME).
@@
 To improve the predictability and actionability of the results for doctors, the clinical factors that are most significant to the prediction are shown on the right [42].
-Reliable medical AI implementation and widespread usage need substantial research and social consensus on prerequisites such as explainability, fairness, privacy protection, and robustness [32].
-Optimized requirements and standards must be satisfied in any therapeutic environment where AI is used, and these requirements and standards need to be updated often.
-Furthermore, laws must be put in place specifying who is in charge in the event of a medical AI-related mishap or accident—designers, researchers, medical personnel, or patients [42].
+For medical AI to work reliably and be widely used, we need to do a lot of research and reach an agreement on important features like explainability, fairness, privacy, and reliability [33].
+We must meet clear requirements and standards in any healthcare setting that uses AI, and we need to update these regularly.
+Additionally, we should establish laws that clarify who is responsible if something goes wrong with a medical AI whether that’s the designers, researchers, healthcare workers, or patients [43].
 3.2 Explainability and Interpretability of Autonomous Systems
-Explainability and interpretability in the context of autonomous systems refer to the ability to understand and make sense of the systems’ decisions and behaviors.
-Explainability refers to an autonomous system’s ability to provide clear arguments for its decisions and behaviors [43].
-It is critical for increasing acceptance and confidence in AI systems, particularly in areas such as banking, healthcare, and autonomous autos.
-While explainability and interpretability are closely connected, interpretability places more emphasis on the capacity to comprehend the internal workings and procedures of the autonomous system [44].
-An interpretable system provides users with insight into the elements and criteria considered while making decisions, allowing them to comprehend how the system came to its findings.
+Explainability and interpretability are crucial concepts in the context of autonomous systems, referring to the ability to understand the decisions and behaviors of these systems.
+Explainability involves an autonomous system’s capacity to provide clear justifications for its actions and choices [44].
+This clarity is essential for fostering acceptance and confidence in AI systems, especially in critical fields such as banking, healthcare, and autonomous vehicles.
+While explainability and interpretability are closely related, interpretability focuses more on understanding the internal mechanisms and processes of the autonomous system [45].
+An interpretable system offers users insight into the factors and criteria that influence its decision-making, enabling them to grasp how the system arrived at its conclusions.
 The research paper The research article [18] focuses on trust and dependability in autonomous systems.
@@
 Following years of research and development, today’s automated technologies represent the peak of progress in computer recognition, responsive systems, user-friendly interface design, and sensing automation.
-According to [43], the global market for automotive intelligent hardware, operations, and innovation is expected to rise from \$1.25 billion in 2017 to \$28.5 billion by 2025.
-According to Intel’s research of the predicted advantages of autonomous cars, employing these advances on public roads would decrease commute time by 250 million hours annually and save over 500,000 lives in the United States alone between 2035 and 2045 [43].
-Modern automobiles employ artificial intelligence (AI) for several tasks, including intelligent cruise control, automatic driving and parking, and blind-spot identification (Figure 111).
+According to [44], the global market for automotive intelligent hardware, operations, and innovation is projected to grow significantly, increasing from $1.25billionin2017to$28.5 billion by 2025.
+Intel’s research on the expected benefits of autonomous vehicles indicates that implementing these technologies on public roads could reduce annual commute times by 250 million hours and save over 500,000 lives in the United States between 2035 and 2045 [44].
+Modern cars utilize artificial intelligence for various functions, including intelligent cruise control, automatic driving and parking, and blind-spot detection (Figure 11).
 Authors [18] describe the challenges of autonomous systems, like, people sometimes tend to be overly excited about the potential of new ideas and ignore, or at least appear to be unaware of, the potential drawbacks of cutting-edge developments.
-Even in the early stages of robotics and autonomous system implementation, humanity preferred to put up with faulty goods and services, but they have gradually come to understand the importance of trustworthy and dependable autonomous systems [zablocki2022explainability].
+Even in the early stages of robotics and autonomous system implementation, humanity preferred to put up with faulty goods and services, but they have gradually come to understand the importance of trustworthy and dependable autonomous systems.
 Numerous examples have demonstrated how operators’ use of automation is greatly impacted by trustworthiness.
-Figure 11: An automated vehicle that provides a valid and understandable rationale for the decision it made at that particular instant, acting as the archetypal example of XAI in automated driving [18].
-When artificial intelligence (AI) has become prevalent in autonomous vehicle (AV) operations, user trust has been identified as a major issue that is essential to the success of these operations.
-Explainable artificial intelligence (XAI), which calls for the AI system to give the user explanations for every decision it makes, is a viable approach to fostering user trust for such integrated AI-based driving systems [47].
+Figure 11: An automated vehicle that provides a clear and understandable rationale for its decisions at that moment serves as a prime example of explainable AI in automated driving [18].
+When AI has become prevalent in autonomous vehicle (AV) operations, user trust has been identified as a major issue that is essential to the success of these operations.
+XAI, which calls for the AI system to give the user explanations for every decision it makes, is a viable approach to fostering user trust for such integrated AI-based driving systems [48].
 This work develops explainable Deep Learning (DL) models to improve trustworthiness in autonomous driving systems, driven by the need to improve user trust and the potential of innovative XAI technology in addressing such requirements.
 The main concept of this [48] research is to frame the decision-making process of autonomous vehicles (AVs) as an image-captioning task, generating textual descriptions of driving scenarios to serve as understandable explanations for humans.
-The proposed multi-modal deep learning architecture (Shown in Figure 12), based on Transformers, effectively models the correlation between images and language, generating meaningful descriptions and driving actions.
-Its contributions lie in formulating the traditional AV decision process for explainability, developing a fully Transformer-based model for generating descriptions and actions, and demonstrating superior performance over baseline models.
-The outcome is a model that enhances user trust, provides insights for AV developers, and offers superior interpretability through its attention mechanisms and end-to-end goal induction.
+The proposed multi-modal deep learning architecture, shown in Figure 12, utilizes Transformers to model the relationship between images and language, generating meaningful descriptions and driving actions.
+Key contributions include improving the AV decision-making process for better explainability, developing a fully Transformer-based model, and outperforming baseline models.
+This results in enhanced user trust, valuable insights for AV developers, and improved interpretability through attention mechanisms and goal induction.
 Figure 12: The Transformers-based multi-modal deep learning architecture that is being suggested [48]
-This research [48] aims to investigate the integration of explainable artificial intelligence (XAI) into autonomous vehicular systems to improve transparency and human trust.
+This research [49] aims to investigate the integration of XAI into autonomous vehicular systems to improve transparency and human trust.
 It delves into the functioning of multiple inner vehicle modules, emphasizing the importance of understanding the vehicle’s decision-making processes for user credibility and reliability.
@@
 AI systems must successfully inform the individuals who share the ecosystem about their objectives, intentions, and findings as the first step toward collaboration.
-In the future, people will work ”with” automation rather than ”around” it, thanks in part to the systematic approach to XAI.
-This research [50] focuses on Explainable Artificial Intelligence (XAI) applications in the process industry.
-The research argues that current AI models are not transparent enough for process industry applications, and highlights the need for XAI models that can be understood by human experts.
+We can hope people to work ”with” automation rather than ”around” it, thanks in part to the systematic approach to XAI.
+This research [51] focuses on XAI applications in the process industry.
+The research argues that current AI models are not transparent enough for process industry applications and highlights the need for XAI models that can be understood by human experts.
 The main contribution is outlining the challenges and research needs for XAI in the process industry.
@@
 Figure 13: Assessing the user’s interaction with XAI [27].
-The precise position of each XAI domain and how they relate to the human user are shown in Figure 13.
-The majority of AI system explanations that are given are usually static and only contain one message [66].
-Understanding cannot be attained by explanations alone [67].
-Because most existing XAI libraries lack user involvement and customization of explanations, users should be able to explore the system using interactive explanations to gain a better understanding of it.
-This is a promising research direction for extending the XAI field [67] and [66].
-To improve human-machine cooperation and move beyond static explanations, a number of efforts have also been proposed.
-Explainable Artificial Intelligence (XAI) has great promise for redefining the relationship between humans and AI systems as it stands at the nexus of technological innovation and societal integration.
-As AI technologies advance, it is more important than ever to ensure accountability and transparency.
-Within this framework, XAI becomes a crucial facilitator, entrusted with shedding light on the murky inner workings of AI models and cultivating user confidence.
-A wide range of breakthroughs are anticipated in XAI, from heightened model transparency and human-centric design principles to regulatory compliance requirements and the rise of hybrid AI systems.
-XAI approaches will place a high value on user-centric design, providing explanations that are both actionable and understandable.
-This will foster acceptance and confidence in AI systems.
-Moreover, it is anticipated that regulatory frameworks would require the incorporation of XAI in essential applications, guaranteeing compliance with accountability and transparency norms.
-Future XAI systems will be distinguished by their contextual sensitivity and interactive explanations, which will enable users to interact with AI decisions in real time and adjust to a variety of situations.
-To guarantee that AI systems follow ethical standards and social values, as well as to democratize access to XAI technologies, efforts must be made to enhance digital literacy and address ethical challenges.
-Fundamentally, XAI’s success is predicated on its capacity to bridge the communication gap between AI systems and human users, fostering mutual respect, trust, and collaboration in an increasingly AI-dependent world.
-This study [68] offers a thorough analysis of Explainable Artificial Intelligence (XAI), focusing on two primary areas of inquiry: general XAI difficulties and research directions, as well as ML life cycle phases-based challenges and research directions.
+Figure 13 illustrates the precise location of each XAI domain and its relationship with the human user.
+According to [67], many explanations of AI systems tend to be static and convey only a single message.
+However, explanations alone do not facilitate true understanding [68].
+To enhance comprehension, users should have the ability to explore the system through interactive explanations, as most existing XAI libraries currently lack options for user engagement and explanation customization.
+This represents a promising avenue for advancing the field of XAI [68] and [67].
+Additionally, various efforts have been made to improve human-machine collaboration by moving beyond static explanations.
+Explainable AI offers a way to improve how people interact with AI systems.
+As AI technology grows, it is crucial to ensure that these systems are accountable and transparent.
+XAI helps by clarifying how AI models work and building trust among users.
+We can expect many new developments in XAI.
+These include making AI models more open, focusing on human-centered designs, ensuring compliance with regulations, and creating hybrid AI systems.
+XAI will prioritize designs that are easy for users to understand and provide clear explanations.
+This clarity will help build trust and encourage more people to use AI systems.
+Regulatory frameworks are likely to require the use of XAI in important areas to ensure accountability and transparency.
+Future XAI systems will need to be sensitive to context and provide interactive explanations.
+This will allow people to engage with AI decisions in real time and adapt to different situations.
+We must also work to improve digital literacy and tackle ethical issues to ensure that AI systems follow moral principles and society’s values, making XAI technologies accessible to everyone.
+The success of XAI depends on its ability to bridge the communication gap between AI systems and human users, which encourages cooperation, mutual respect, and trust in an increasingly AI-driven world.
+This study [69] offers a thorough analysis of XAI, focusing on two primary areas of inquiry: general XAI difficulties and research directions, as well as ML life cycle phase-based challenges and research directions.
 In order to shed light on the significance of formalism, customization of explanations, encouraging reliable AI, interdisciplinary partnerships, interpretability-performance trade-offs, and other topics, the study synthesizes important points from the body of existing literature.
@@
 5 Conclusions
-Explainable Artificial Intelligence (XAI) is gaining popularity in a range of fields due to its critical role in addressing critical issues connected to AI adoption.
-As AI systems become more integrated into society, transparency and interpretability become increasingly important.
-By offering tools to clarify how AI models make decisions, XAI helps users develop a sense of confidence and comprehension.
-XAI’s primary objective is to make AI models clear and intelligible.
-With the help of XAI, the general public will be able to peer inside the black box and comprehend the aspects that affect the AI’s decision-making process.
-The paper discusses the essential details of XAI and offers a comprehensive overview for a solid understanding.
-Furthermore, this article discusses in detail the three main application fields of XAI.
-Lastly, the authors attempt to outline the difficulties in applying XAI and suggest potential future paths.
+XAI, or Explainable Artificial Intelligence, is becoming important in many industries because it helps solve key challenges with using AI.
+As AI becomes more common in our daily lives, understanding how it works is essential.
+XAI provides tools that help people see and understand how AI models make decisions.
+The main goal of XAI is to make these models easier to understand.
+It allows people to look inside the ”black box” of AI and see what affects its decisions.
+The paper gives a clear overview of the key parts of XAI.
+It also discusses three main areas where XAI can be applied.
+Finally, the authors talk about the challenges of using XAI and suggest possible future directions.
 […]